### 14 Jan 2024 v1

# Privacy-Preserving Intrusion Detection in Software-defined VANET using Federated Learning with BERT

Abstract
The absence of robust security protocols renders the VANET (Vehicle ad-hoc Networks) network open to cyber threats by compromising passengers and road safety. Intrusion Detection Systems (IDS) are widely employed to detect network security threats. With vehicles’ high mobility on the road and diverse environments, VANETs devise ever-changing network topologies, lack privacy and security, and have limited bandwidth efficiency. The absence of privacy precautions, End-to-End Encryption methods, and Local Data Processing systems in VANET also present many privacy and security difficulties. So, assessing whether a novel real-time processing IDS approach can be utilized for this emerging technology is crucial. The present study introduces a novel approach for intrusion detection by the capabilities of Federated Learning (FL) in conjunction with the BERT (Bidirectional Encoder Representations from Transformers) model for sequence classification (FL-BERT). The significance of data privacy is duly recognized. According to FL methodology, each client has its own local model and dataset. They train their models locally and then send the model’s weights (or parameters) to the server. After aggregation, the server aggregates the weights from all clients to update a global model. After aggregation, the global model’s weights are shared with the clients. This practice guarantees the secure storage of sensitive raw data on individual clients’ devices, effectively protecting privacy. After conducting the federated learning procedure, we assessed our models’ performance using a separate test dataset. The FL-BERT technique has yielded promising results, opening avenues for further investigation in this particular area of research. Also examined were several conventional machine learning models such as Random Forest (RF), Support Vector Machine (SVM), Logistic Regression (LR), and K Nearest Neighbours (KNN) to detect threats. However, these models lack data privacy. We used the Vehicular Reference Misbehavior (VeReMi) benchmark dataset in both cases. We reached the result of our approaches by comparing existing research works and found that FL-BERT is more effective for privacy and security concerns. Overall, our results suggest that FL-BERT is a promising technique for enhancing attack detection effectiveness, whereas the FL-BERT approach performs satisfactorily in attack detection and safeguarding data.
keywords:
FL-BERT, VeReMi, SDN, VANET, IDS, Cybersecurity Privacy
Introduction
VANETs (Vehicle ad-hoc Networks) are a subclass of Mobile Ad-hoc Networks, often described as MANETs. Rapid topology changes influence a specific type of wireless multi-hop network called a VANET because of the high node mobility. Inter-vehicle communication is becoming potential research, standardization, and development as more cars are outfitted with wireless communication and computing technology. The new era of automobiles and road traffic systems will depend on wireless networks and machine learning. As stated previously, many lives have been lost on highways. Consequently, when road traffic increases, road safety must also grow. We need a mechanism by which vehicles can be made sufficiently intelligent to manage their process for road safety. This notion constituted the foundation for VANETs. VANETs enable vehicles to speak with one another and learn whether a diversion is coming or if more data is necessary. Therefore, a VANET based on IEEE 802.11p facilitates communication between vehicles. SDN-based VANETs follow the below architecture [15]. VANET is a wireless network in which both stationary and mobile nodes connect. Such a system has frequent disconnections, a dynamic topology, and significant node mobility. Each attribute may result in a security flaw that hazards the system. In comparison to MANET, VANET mobility is considerably greater. In a VANET, vehicles travel arbitrarily throughout the network and are restricted by the network’s topology. To detect known and unencrypted attacks in a local detection scenario, researchers are showing a great interest in current years due to the SDN concept in VANET dealing implementation of the learning-based detection mechanism. IDS is still in its infancy, and many studies must be conducted to make IDS as successful as possible. Current IDS is plagued by difficulties and obstacles that require rapid and intensive investigation. This work has identified significant elements and activities that research communities must address.
Several basic study gaps in learning-based IDS research have been noted in the literature review below. Still, there is enough opportunity to improve the Intrusion detection Model accuracy with data privacy and security concerns. Moreover, IDS users still struggle to detect known unencrypted attacks because IDS has not yet achieved 100% accuracy. One significant issue with IDS is that they regularly alert the auditor to false positives. However, an IDS can be optimized to reduce the number of false positives and false negatives. Robust and Realistic IDS Model: In VANETs and IDS, adversarial attacks can impact or mislead machine learning models by purposely feeding them erroneous or malicious data. This attack might be tailored to target and exploit the weaknesses of a particular IDS model precisely. Deficiency of publicly available VANET database with attacks. No fast, inexpensive, and effective in-built privacy-preserving machine learning-based IDS is still visible.
Vehicular ad hoc networks (VANETs) demand frequent communication between cars, transmitting data that may compromise sensitive information. The challenge is balancing the ongoing exchange required for optimal performance and safeguarding user privacy. The constant transmission of position data poses a vital privacy issue. The patterns of an individual’s daily travel or frequently visited places possess significant value, not just for commercial purposes but also for possibly nefarious commitments. Moreover, in light of the increasing prevalence of open-access situations and the rise of devices engaged in communication within the VANET environment, the network offers a favorable environment for possible unauthorized access and spying. Nowadays, Developing an IDS system that facilitates continuous monitoring with high attack detection accuracy while safeguarding against unauthorized access or unintended use of data is a more significant task for researchers to find solutions.
In consequence, the contributions of this paper are as follows:
  1. 1.
    This paper proposes an innovative approach called FL-BERT to address privacy and security issues in the Software-defined VANET that combines the capabilities of Federated Learning (FL) with the BERT model to enhance sequence classification tasks.
  2. 2.
    In addition, conventional machine learning models were trained, particularly Random Forest (RF), Support Vector Machine (SVM), Logistic Regression (LR), and K-Nearest Neighbours (KNN) to compare evaluation reports with FL-BERT to measure effectiveness in terms of attack detection. To demonstrate that FL-BERT can perform better despite its complex structure to maintain data privacy without compromising the model’s performance.
  3. 3.
    The approaches employed in our study have been evaluated using the VeReMi dataset, characterized by its substantial size and comprehensive level of information. Utilizing this dataset enables us to enhance our comprehension of the efficacy of our security techniques in detecting diverse security risks and employing privacy. This comprehensive analysis of our methodologies allows us to identify our approach’s merits while highlighting potential areas for enhancement.
  4. 4.
    Our research also discovers the contextual capabilities of BERT significantly improve threat detection, outperforming conventional models in terms of adaptability and performance.
Related Works
As a part of the research, different articles have been reviewed to explain how earlier researchers have ended work on an Intrusion Detection System for cyber-attack detection in SDN-based VANET. This Literature review study employs the methods stated by PRISMA. The methodology covers 8 main steps: User needs assessment, define the Research Questions, Search String, Selection of Source, Select Relevant Paper, Screening Findings, Extract Data, and finally, synthesize the collected data. (Kitchenham and Charters, 2007) To Discover the user needs, the following steps are essential to conclude What users want or need to protect themselves from cyber-attacks. Government and private organizations are the two main stakeholders in the cybersecurity field. In the private sector, companies from various sectors of the economy, the ISPs, software vendors, end users, etc.
Table 1: VeReMi Dataset attacks and size.
Attacks Size
BENIGN 437429
Attack type 16 (Eventual stop Attack) 56595
Attack type 4 (Random Attack) 30510
Attack type 2 (Constant Offset Attack) 30473
Attack type 1 (Constant Attack) 30473
Attack type 8 (Random Offset Attack) 29460
IDS and Malware detection systems can still not detect unknown, Encrypted, and Inference attacks with 100% accuracy. According to Kaspersky, most Malware detection products do not even reach the required 90% security threshold. A significant player in the IDS market, CISCO, claimed that Rapid detection—those made in less than five minutes—and long-term learning capacities displayed excellent performance in their IDS. Cisco Encrypted Traffic Analytics, or IDS discovered nearly two-thirds of all known malicious flows in less than five minutes. Moreover, Cisco declared this IDS demonstrated better identification results even with low volumes of 0 to 20 flows, whereas 2000 plus flows enabled 100% detection for known attacks. Nevertheless, this solution has a limitation with a small attack size and taking three hours to gain 100% accuracy. However, Cisco ASDM to manage Cisco Secure Firewall ASA (with IDS/IPS) has been impacted by vulnerabilities CVE-2022-20829, CVE-2021-1585, and CVE-2022-20828. Researcher Jake Baines, who disclosed them to Cisco in February and March of 2022, added that many users were undoubtedly not updating their Cisco firewalls properly. [9] [21].
There are several difficulties in creating an adaptable and versatile IDS using ML in SDN-based VANET networks due to the need for effective data classification in academic research. To provide a consistent and precise evaluation of intrusion detection systems, robust datasets must be created. Several datasets are available for intrusion detection testing and evaluation. For example, a benchmark dataset for misbehavior detection systems for VANETs is the VeReMi dataset. It is a synthetic dataset made with LuST (Version 2) and VEINS (with modifications based on Version 4.6). In this dataset, onboard unit message logs from a simulation environment are included, along with tagged ground truth. It has two primary uses: it gives a starting point for determining how well misbehavior detection systems work city-wide. Though this database has very few attack classes and sizes. [13] [23]. Many researchers are generating synthetic datasets for their research purposes by using different simulators, but their database size is shallow; the entry size is mentioned in the table.
The article discusses the ML-based tactic to detect DDoS Attacks in SDVN. [26]. This paper suggests a method for spotting threats in SDN-based vehicular networks based on machine learning. The research has a particular emphasis on vehicle-to-infrastructure (V2I) links. It analyzes the current environment to establish the machine learning strategy that will most successfully detect threats posed by distributed denial of service (DDoS) attacks (Short-lived TCP Flooding, UDP Flooding). To accomplish this, it collects seven different elements of actual data from TCP and UDP flows both when DDoS attacks them and when they are not. The network is educated by employing several different supervised learning algorithms depending on the data acquired, and it is discovered that the gradient boost classifier produces the most satisfactory outcomes. In addition, the random forest, decision tree, and linear SVM approach all produce results that are pretty close to being the best. On the other hand, the limited training data could be to blame for the NN’s weak performance. Accurate, measurable data was not found in the write-up or the chart.
[5] presents an AI-based system to identify a Route Request (RREQ) flooding attack. Researchers used a genetic algorithm to optimize feature subset selection and an ANN for classification. The model scored 99% accuracy from 1000 entries simulated dataset.
[30] In this study, they analyze DDoS assaults in an SDN-based vehicular network by employing TCP flood, UDP flood, or ICMP (internet control message protocol) flood. In this research paper, the authors present a scheme for detecting Distributed Denial-of-Service (DDoS) attacks in SDVN atmospheres using three separate models. A “trigger detection model” for inbound packets and another method is the “flow table feature-based detection model.” Both methods use the OpenFlow protocol features and the detection of a threat model based on support vector machine classification. All of these models are used in conjunction with one another. They synthesized and simulated network traffic to accommodate various DDoS attacks via Scapy and hping3, and the findings suggested an efficiency of over 97%. In the article, ML-Based Intrusion Detection Method for VANETs.
[31] the authors developed an ML-based Senior2Local IDS with propagation to identify potential threats in VANETs both Globally and Locally. They used a fuzzification ANN technique to recognize malicious attacks globally to protect CHs in RSUs. After that, a game theory trust system is developed separately or RSUs. They employed a simple SVM to detect attacks locally. The findings demonstrated that the method was more reliable and resilient than earlier ML-based tactics.
[10] conduct the experiment on a ToN-IoT dataset and test various ML approaches in binary and multiclass classification problems in VANET to detect intrusion. Using LR, binary classification problems can be addressed. To handle multiclass challenges, researchers used a one-vs-rest multiclass strategy.
[17] suggests employing data mining methods in VANETs to identify existing assaults and find other attacks that have not yet been discovered. The suggested approach has three primary contributions:
(A) A decentralized vehicle network and build scalable communication.
(B) The use of two data mining models to prove viability for an IDS in VANETs;
(C) The discovery of novel patterns of unidentified threats.
The suggested solution uses a cell grid to organize the previously mentioned network. Every cell has a transmission tower, enabling users to communicate with others and the internet. Each will carry out its data mining models and rules to identify new assaults. Therefore, the IDS can develop new rules that may be communicated for each subnetwork. The trucks and the tower cell collect the data sent throughout the network. The authors use Naive Bayes and Logistic Regression classifiers to analyze the obtained data. Finally, the authors use simulation to evaluate their strategy’s performance over a network. First, the SUMO simulator produces a file that tracks movement. Then, to mimic a wireless network, this information is given into ns-2. This scenario includes 150 cars that turn randomly at junctions, drive at the posted speed limit (which ranges from 5 to 20 meters per second), and stop at traffic lights that are randomly positioned. To validate the IDS, 5 automobiles were outfitted with Linux and made to run various network applications. After that, nine months were spent using TCPdump to generate the dataset. There were 4 different kinds of attacks, with 39 different types documented. After that, WEKA is used to categorize them. Finally, the metrics recall, F-measure, and Matthews’s correlation coefficient were utilized to assess the models (MCC).
[28], With the help of feature selection and hyperparameter optimization, the DT classifier attained an accuracy score of 99.35%, which was the best possible result. The excellent results achieved with the methodology developed proved the considerable significance of feature selection and optimization techniques. These findings were proven by the outstanding results obtained with the method.
[14] proves the existence of a collaborative security threat detection method within software-defined vehicular cloud architecture. A traditional vehicular environment has exceptional features, such as high flexibility, dynamic networking, and rapid connectivity. His proposed model said an SDVC controller had collected the flow information from the vehicular cloud (VC) to train multiclass SVM. In the next step, the SVM classifier identifies the bad actors among the network’s nodes. However, performance is still under 90% for Accuracy score, Precision, and Recall.
However, [19] offers an LA-based system for IDS in VANET. The suggested methodology protects users’ anonymity by giving each car a unique, changing ID. The suggested approach involves developing a VANET administration system. There is a base station at the beginning and end of each route, and all cars have transmitters to stay in touch with the base stations. To redirect traffic, each attacker will generate malicious packets. The budgets of the attackers and the system are similar. The more the attacker’s resources, the greater the number of packets it can create. All that is needed to identify the attackers is their dynamically assigned identifiers. If unique IDs are assigned to vehicles, and many vehicles have the same ID, then at least one of the vehicles is likely malevolent. Under attack, grid budgets may be attributed, and sampling rates can be recalculated with the help of learning automata. Simulations run on their simulator are used to evaluate models. The success rate for detecting four attackers ranges from around 44% to 77%. In the suggested setup, the IDS is located in each base station and has complete sight of overall network traffic.
[27] proposes a BUSNet-based IDS tailored to protect VANETs from malicious activity. This hierarchical anomaly detection technique is built on the virtual setup of the mobile bus agents. In this configuration, the buses serve as the cluster heads, collecting and relaying data packets broadcast by passing vehicles through accessing points installed along the roadways. After that, an NN-based classification system takes that data and uses it to spot DDoS assaults. The proposed method is tested with the ns-2 network simulator in an artificial network. The researchers do not disclose the method or traffic simulator utilized to generate the simulation results. Fifty unique vehicles are simulated, each communicating using a constant bit rate (CBR) and 512-byte packet size. During each attack, two vehicles work together to send out updates at small durations of 0.01 seconds at four different checkpoints over 10 seconds. Both vehicles are used in every attack. The authors establish their benchmark by picking a value below which they treat everything as an assault. This number might range from 0.05 to 0.7, which helps to explain the discrepancy in these results. A threshold value of 0.2 is recommended as the best option. Overall, accuracy has not achieved the milestone to protect VANET.
[18], The proposed fix is a detection system that uses a statistical method, linear regression, and a watchdog to discover greedy behavior in the MAC layer passively. A watchdog monitors the degree to which active nodes’ access periods correlate as part of the suggested approach. The network is under assault if the correlation coefficient is close to 1 but not precisely 1, and the linear regression straight-line slope is also not close to 1. This can happen when the correlation coefficient is near 1 but not quite 1 or when the slope is very close to 1 but not quite 1. The developed program does this by monitoring three metrics: the interval between transmissions, the total time spent transferring data, and the number of connection attempts made by a node. Then, with the aid of the trace file created by SUMO, the solutions’ functioning is analyzed in real time. -3. The SUMO scenario is based on a city map with labels and signals. Researchers first simulate the network using nodes that behave correctly to ensure the linear regression method is reliable. After that, four greedy automobiles are produced by injecting greedy nodes one at a time. To detect greedy behavior, this approach takes 1.3 seconds for 1 node, 1.9 seconds for 2 nodes, 3.1 seconds for 3 nodes, and 7.9 seconds for 4 nodes. [2] The solution proposed in the paper is an ANN-based ill-use and anomaly intrusion detection system that can identify DoS assaults. The authors designed the solution by first developing a mobility scenario within SUMO. For ns-2 to read the files SUMO created, a conversion was performed using MOVE. Ultimately, ns-2 was utilized to imitate the vehicles’ communication ability. When creating the mobility and traffic situation, the authors took the urban mobility of Manhattan as their inspiration. The simulation takes two minutes and fifty seconds with thirty cars and six Road Side Units (RSUs). A CBR (Constant Bit Rate) traffic that conveys User Datagram Protocol (UDP) packets is being performed by the vehicles. Only one of the automobiles intentionally tries to cause harm. It will discard the packets rather than send them to the destination. The solution built was accurate 85.02% of the time when classifying normal behaviors and 98.45% when classifying deviant behaviors.
[12] The authors advocate using a POS approach to reduce the number of features taken from the trace file. These are fed into an artificial neural network to teach it to classify data. After that, the ns-2 trace files that were created were employed in the POS algorithms so that the appropriate characteristics could be extracted. Next, Fuzzification is applied to the dataset’s data to prevent classification difficulties. Finally, the dataset is classified with the help of an FFNN using the IDS created for this solution. Sixty thousand records from the dataset were used, with half designated for training, another quarter for testing, and the remaining quarter for validation (25%). This dataset includes both benign and harmful activity in equal measure. The malicious behavior that was designed was a Black Hole attack, in which the malicious vehicle would not forward any received packets but dump them instead. Ultimately, the intrusion detection system was evaluated using anomaly and misuse detection. Regarding identifying misuse, it acquired a classification of 99.89 % for normal conduct and 99.80% for deviant behavior. Regarding detecting anomalies, the findings for typical and aberrant behaviors came in at 99.87 and 99.72%, respectively.
[22] This paper presents a cluster-based intrusion detection system (IDS) to defend a network against selective forwarding, black holes, wormholes, packet repetition, resource depletion, and Sybil attacks. The technique that has been presented uses several different detection managers that operate on three different levels: cluster associate, cluster head, and RSU. The cars’ speeds determine how they should be organized into clusters. Cluster connection and the assurance of data safety are the criteria utilized in selecting the cluster heads. When social behavior is introduced, the connectivity inside a cluster is improved. The IDS architecture comprises two primary detection schemes and a decision scheme. LIDS and GIDS are the detection systems run at cluster members and heads. The name of the decision-making scheme is Global Decision System, which is hosted on the RSUs. Because of this, the system can monitor several organizations and identify threats on multiple levels. In addition, each level can perform a unique set of algorithms and detecting approaches, which evaluate a unique set of characteristics. The Global Decision System will be provided with the collective standing of each car, which the cluster head will then send. The Worldwide Decision Scheme will then calculate the trust level of every vehicle. To evaluate the proposed system, ns-3 was used as the network simulator, while SUMO was used to mimic the vehicle’s movement. The IDS was evaluated concerning the detection rate, the false positive rate, and the detection time. The various assaults were put through their paces using a range of vehicle counts during testing. The percentage of hostile cars always equals 45% of the total. The provided findings range from 92% to 100%, although this depends on the total number of cars and the type of attack conducted.
Table 2: Review of different ML-based IDS models in VANET and SDVN.
Reference ML Models Dataset Attack Featuring Method Top Result
[26]
GB, RF, DT,
Linear SVM,
LR, NN,
Neural Net.
Generated
Short-lived
TCP Flood,
UDP Flood.
Not measurable
from the chart
[5]
ANN
NS2 Trace
RREQ Flood GA based
Accuracy: 95%,
Precision: 97%,
F-Score: 98%,
FPR: 3%
[30]
SVM
DARPA 2000,
CAIDA
DDoS 2007
TCP flood,
UDP flood,
ICMP Flood,
Etc.
Flow-based
features
and statistical
based entropy
TCP Traffic
Pre-Selection:
DR 98.26 %,
FR 0.52%
After-Selection:
DR 98.56%
FR 0.32%
[32]
DeepVCM,
KNN, DT,
1D-CNN, LSTM
ISCX2012,
NS3 Trace
Brute Force,
SSH DDoS,
HttpDoS
RAW data
NS3 Dataset/
DeepVCM/DoS
Precision 98.5%
Recall 98.3%
F1-Score 98.4%
[31]
ANN Generate -
CNN
algorithm
applied to
Extract features.
Accuracy:
SMV-CASE 98.7%,
CEAP 98.9%,
Senior2Local 98.37%
[10]
LR, NB,
DT, SVM,
kNN, RF,
AdaBoost,
XGBoost
ToN-IoT
DoS
DDoS
etc.
Features
Selection:
a. Apply Chi²
b. Apply
SMOTE
c. Apply above
both
With All Features
XGBoost:
Accuracy 99.1%,
Precision 98.4%,
Recall 99.1%,
F1-Score 98.7%
[28]
kNN, SVM
DT
Generated DDoS
MRMR
filter
method
DT Algorithm with
25 Features:
Accuracy 99.35%,
Sensitivity 99.22%,
Specificity 99.80%,
F1-Score 99.21%
[14]
SVM KDD CUP 1999
DoS, Probing,
U2R, R2L
Etc.
Under Diff. Attack
(Approx. from
Graph when $\alpha$=0):
Accuracy
SVM-VC 68.00%,
SVM Nearest
Neighbour 57.00%,
SVM Individual 39.50%
[17]
NB, LR TCPdump
DoS, R2L,
U2R, Probing
Not measurable
from Chart.
LR showed the
best result.
[2]
ANN NS2 Trace file DoS - Accuracy 85.02%
[12]
ANN
NS2 Trace file and
Animator (32000)
Black Hole - Accuracy 85.02%
[1]
FFNN,
SVM
NS2 Trace file
Grey Hole,
Rushing
Accuracy:
SVM-Abnormal
99.80%
[24]
SVM NS2 Trace file
Wormhole,
Selective Forwarding,
Packet Drop
Hybrid
Fuzzy
Multi-
Criteria
Feature
Selection
Performance
evaluated
by DR, FPR, DT
[16]
Learning
automata
Generated Flooding, Blackhole - -
[25]
MLP (ANN),
BiGAN
KDD99,
NSL-KDD
DoS, U2R,
R2L, Probing etc.
Accuracy 98.4%
Precision 95.23%
Recall 96.74%
F1-score 95.98%
Precision and recall
are low due to
complex model.
[6]
Hybrid model
based on RF
KDD99,
CICIDS2017
DoS, DDoS,
HeartBleed Etc.
- Accuracy 96.93%
[20]
Softmax
Generate
DDoS - Accuracy 96.9%
[4]
SVM, GA,
PSO, ACO
NSL-KDD
DoS, Probing,
Unauthorized
access
- Accuracy 98%
[11]
RF
NSL-KDD,
UNSW-NB15
DDoS - Accuracy 99.95%
[3]
CNN-LSTM,
CNN- bidirectional
LSTM, LSTM,
Auto-encoders
Stacked LSTM,
GRU
VeReMi
Various types
of Anomalies
- Accuracy 98%
Figure 1: Overall FL-BERT model diagram.
Methodology
In this section, we describe the selected datasets and attacks, as well as the implementation of our proposed FL-BERT model alongside conventional models, including Random Forest (RF), Support Vector Machine (SVM), Logistic Regression (LR), and K-Nearest Neighbours (KNN). All of these are intended to improve the detection of anomalies in Software-defined VANET architecture. In addition, the FL-BERT data privacy mechanisms inherent to this architecture are discussed.
The VeReMi dataset was created to test the effectiveness of VANET misbehavior detection systems (vehicular networks). This database contains onboard unit message logs created from a simulation environment and tagged ground truth. The collection contains malicious messages that are meant to cause erroneous application behavior, which misbehavior detection techniques are designed to stop. The initial dataset includes five types of position falsification attacks. [13] Database source: This data was generated by [13], and It has been collected from [23].
To achieve the best possible hardware support for running the code, Microsoft Azure ML Studio VM instance with Python 3.8, PyTorch 1.13.1, TensorFlow 2.13.0, Protobuf 4.24.3, and Transformers 4.33.2.
Random Forest Classifier
The randomization must decrease correlation while retaining strength to increase accuracy. The forests under study grow each tree by applying inputs or combinations randomly chosen at each node. Accuracy results from the produced forests are comparable to Adaboost favorably. This class of methods has the following desirable qualities:
A. Its accuracy is comparable to Adaboost and occasionally much better.
B. It is relatively resistant to noise and outliers.
C. It moves more quickly than bagging or boosting.
D. It is straightforward, easily paralleled, and provides helpful core estimates of error, strength, correlation, and variable relevance. [8]
Logistic Regression
LR is a linear model used to classify instead of predict. A logistic function is used for probabilities that describe the possible results of a single trial in this model. Typically, a sigmoid curve with the shape of a Sigmoid curve is called a logistic function or logistic curve.
$\boldsymbol{f}\left(\boldsymbol{x}\right)\boldsymbol{=}\frac{\boldsymbol{L}}{\boldsymbol{1}\boldsymbol{+}{\boldsymbol{e}}^{\boldsymbol{-}\boldsymbol{k}\left(\boldsymbol{x}\boldsymbol{-}\boldsymbol{x}\boldsymbol{0}\right)}}$
Equation 1: Logistic Function.
Where x0 is the x value of the sigmoid’s midpoint; L is the curve’s maximum value; k is the curve’s logistic growth rate or steepness. [29]
Support Vector Machines
Support Vector Machines (SVMs) are algorithms used in learning for classification, regression, and outlier detection tasks. They are particularly effective in situations with data and can handle cases where the number of dimensions exceeds the number of samples. SVMs use a subset of training data called support vectors, which helps improve memory efficiency. One advantage of SVMs is their versatility; they can utilize kernel functions, including ones you can define yourself. However, there are some limitations to consider. Overfitting may occur if the number of features greatly exceeds the number of samples. Careful choices must be made regarding kernel selection and regularization techniques to address this issue. Additionally, SVMs do not provide probability estimates by default; obtaining these requires performing five-fold cross-validation. In the scikit learn library, SVMs can handle both sparse vectors; however, when making predictions on data, it is important first to fit the model on similar data. The decision function for an SVM is expressed as: [10]
$f(x)=\sum_{i=1}^{N}\alpha_{i}y_{i}K(x_{i},x)+b$ (1)
Equation 2: Support Vector Machines Function.
Where $\alpha_{i}$ are the Lagrange multipliers, $y_{i}$ are the labels, $K$ is the kernel function, and $b$ is the bias.
K-Nearest Neighbors
The K-Nearest Neighbors (kNN) set of rules, a case primarily based on gaining knowledge of technique, keeps all training facts for class. Its ”lazy” nature limits its use in huge-scale programs like dynamic internet mining. To decorate efficiency, consultant statistics factors can be used to symbolize the whole schooling set, basically developing an inductive studying model. While numerous algorithms like selection trees and neural networks serve this cause, kNN’s simplicity and effectiveness, especially obvious in textual content categorization obligations like the Reuters corpus, encourage efforts to optimize its performance without compromising accuracy. During model production, each information factor possesses a maximal local neighborhood with statistics points of the same elegance label. The largest of these neighborhoods, termed the ”largest global neighbourhood,” acts as a representative. This method repeats till all statistics points are represented. Unlike traditional kNN, this approach would not require a predefined ( k ); it is routinely decided through model production. Using representatives no longer simply reduces records but additionally boosts performance, addressing kNN’s inherent barriers.
This study will address multiclass problems using the One-vs-All or OvA multiclass methodology. The steps for data preprocessing before feeding to conventional models and how to handle the FL-BERT model are discussed below.
Data Preprocess
A. Dataset: VeReMi
A benchmark dataset for misbehavior detection systems for VANETs is the VeReMi dataset. It is a synthetic dataset made with LuST (Version 2) and VEINS (with modifications based on Version 4.6). In this dataset, onboard unit message logs from a simulation environment are included, along with tagged ground truth. It has two primary uses: it gives a starting point for determining how well misbehavior detection systems work city-wide. This database has very few attack classes and sizes, as shown in Table. [13] [23]. As shown in Table, we create a balanced subset of preferred samples across all attack types for model training and evaluation.
Table 3: VeReMi Dataset. (Resampled)
Attacks Size
BENIGN 50000
Attack type 16 (Eventual stop Attack) 28832
Attack type 4 (Random Attack) 30510
Attack type 2 (Constant Offset Attack) 30473
Attack type 1 (Constant Attack) 30473
Attack type 8 (Random Offset Attack) 29460
B. Data Preprocessing
We created a suitable textual feature by linking several columns from the dataset. Additionally, we encode categorical labels utilising the LabelEncoder class. The data is then split into training and testing sets, forming the basis for subsequent model development.
C. FL-BERT Model
The model being discussed in this paper is a practical application of federated learning, specifically utilising the BERT (Bidirectional Encoder Representations from Transformers) architecture. This model has been specifically tailored for use with the VeReMi dataset. Federated learning is an approach in machine learning where the training process is distributed over several devices or servers, referred to as clients, without the need to share the raw data. This approach guarantees the confidentiality of records, as the most efficient iterations of changes are transmitted to a central server, which consolidates these updates to enhance a global model.
Data preparation plays a key role in the pipeline of every machine learning system. In the present model, a combination of various columns extracted from the dataset is performed, including ’sendtime’, ’sender’, ’messageID’, ’pos’, ’spd’, ’sendtime’, and ’sender’, resulting in the creation of a unified column referred to as ’textual content’. This compiled text serves as the input for the BERT model. The column labeled ’AttackerType’, which potentially contains explicit data indicating unique categories of attackers is subsequently encoded using label encoding to convert it into a numerical format. The column labeled ”labels” in this encoding acts as the dependent variable for the model. The data is subsequently divided into educational and testing components, with 80% of the data allocated for educational purposes and the remaining 20% set aside for testing.
Tokenization is a subsequent procedure that entails transforming the factual content of the text into a structured format suitable for input into the BERT model. The utilization of the BERT tokenizer, namely the ”bert-base-uncased” paradigm, serves the intended objective. To facilitate the tokenization process and prepare the records for training, a class called ‘CustomDataset‘ has been implemented. The brilliance of this approach lies in its utilization of texts, related labels, a tokenizer, and a maximum sequence time as input parameters.
In the federated learning framework context, the model encompasses two primary components, namely the ”Client” and the ”Server”. Every buyer corresponds to a device or server with a fraction of the training data. The user initiates a BERT model for classifying collections, transfers it to the designated computing device (GPU or CPU), and specifies an optimizer for the training process. The global model (the model aggregating updates from all local models) is continuously reset and transmitted to all clients. This means that the global model does not excessively rely on data from any specific user, which could potentially disclose patterns specific to that user. By continuously resetting the global model, it is harder for an attacker to infer specific patterns from the actions of the model, as a result enhancing privacy.
On the other hand, the server maintains a global model that consolidates the updates received from all clients. The server magnificence employs the ‘mixture‘ strategy, which involves receiving models from all clients, calculating their weights’ average, and updating the global model with this computed average. This aggregation process ensures that the global model benefits from analysing individual customer models.
The federated training approach entails the process of training models on individual clients, aggregating their updates on a central server, and subsequently distributing the new global model back to the clients. The aforementioned procedure is iterated for a specific number of iterations.
The subsequent section presents an assessment function utilised to evaluate the comprehensive performance of the BERT models. This attribute sets the model to evaluation mode, which generates predictions on the test dataset and compares them with the actual labels. The performance is subsequently evaluated using a classification report, which provides metrics such as precision, recall, and F1-score for each class inside the ’AttackerType’ column.
Ultimately, the proposed version presents a thorough implementation of federated learning using the VeReMi dataset’s BERT architecture. This version provides a robust solution for training on distributed data while ensuring data privacy, by utilising the capabilities of BERT and the privacy-preserving characteristics of federated learning.
D. Conventional Model
In the course of our study, we performed an assessment utilizing machine learning classifiers such as Random Forest (RF), Support Vector Machine (SVM), Logistic Regression (LR), and K Nearest Neighbours (KNN). The primary objective of our study was to detect abnormalities in the ’AttackerType’ within a Software Defined VANET architecture. The dataset was partitioned into training and test subsets, with the ’text’ column serving as the feature.
The Term Frequency Inverse Document Frequency (TF IDF) vectorization technique was employed to preprocess the data for machine learning. The proposed methodology quantifies the significance of individual terms within the dataset by considering their frequency across all documents, resulting in a normalised representation. We have restricted the features to the highest-ranking one thousand to enhance efficiency and prioritise specific terminologies.
We trained each classifier using the modified training dataset following the vectorization process. Subsequently, an assessment was conducted to evaluate their performance on the test dataset. By employing this methodology, we evaluated each classifier’s efficacy and constraints in identifying anomalies. The study of the findings includes precision, recall, and F1 score measures, which identify the most effective models for our specific application case.
Results and Evaluation
In this study, we performed a comparative analysis of machine learning models and the innovative Federated Learning BERT (FL BERT) approach to identify abnormalities in Software-Defined Vehicular Ad Hoc Network (VANET) infrastructures. We evaluated many models, including Random Forest (RF), Support Vector Machine (SVM), Logistic Regression (LR), and K Nearest Neighbours (KNN).
The FL BERT model exhibited an 84% accuracy rate, substantiating its efficacy in processing context-rich data. The system continuously demonstrated high levels of precision, recall, and F1 score metrics across various attacker types, suggesting its capacity to detect a wide range of anomalies accurately.
Table 4: Accuracy Comparison Models Accuracy Random Forest 49 SVM 59 Logistic Regression 59 KNN 38 FL-BERT 84
Figure 2: Accuracy comparison chart
On the other hand, the Random Forest algorithm demonstrated an accuracy rate of 49%, indicating competence in certain categories while exhibiting deficiencies in others. Support Vector Machine (SVM) and Logistic Regression models demonstrated a comparable accuracy rate of 59%. Although these models showed proficiency in identifying certain types of attackers, their performance was comparatively poorer in other areas. This discrepancy suggests the possibility of overfitting or insufficiency in effectively addressing the intricate nature of the dataset.
Table 5: Performance Metrics for Different Models
Model Type Precision Recall F1-Score Support
FL-BERT 0 0.94 0.95 0.95 9899
1 1.00 1.00 1.00 6158
2 0.65 0.88 0.75 6096
4 0.98 0.74 0.84 6181
8 0.69 0.45 0.55 5899
16 0.79 0.95 0.86 5717
Random Forest 0 0.73 0.79 0.76 9899
1 1.00 1.00 1.00 6158
2 0.35 0.35 0.35 6096
4 0.15 0.14 0.14 6181
8 0.12 0.11 0.12 5899
16 0.35 0.33 0.34 5717
SVM 0 0.76 0.83 0.79 9899
1 1.00 1.00 1.00 6158
2 0.53 0.42 0.47 6037
4 0.32 0.34 0.33 6190
8 0.31 0.33 0.32 5892
16 0.48 0.45 0.46 5784
Logistic Regression 0 0.73 0.82 0.77 9899
1 1.00 1.00 1.00 6158
2 0.50 0.46 0.48 6037
4 0.33 0.37 0.35 6190
8 0.33 0.30 0.31 5892
16 0.48 0.39 0.43 5784
KNN 0 0.65 0.77 0.70 9899
1 0.52 0.58 0.55 6148
2 0.21 0.21 0.21 6037
4 0.13 0.13 0.13 6190
8 0.13 0.11 0.12 5892
16 0.35 0.25 0.30 5784
The KNN model had a lower accuracy rate of 38%, indicating that it may not be well-suited for the given data distribution or the intrinsic difficulties of the VANET system. In summary, while standard machine learning models have demonstrated varying levels of efficacy, FL BERT has emerged as a viable choice for detecting abnormalities in Software-Defined VANETs. In handling various types of assailants, this model exhibits superior accuracy and precision compared to its counterparts. This study highlights the potential of utilizing advanced neural network models to address complex cybersecurity concerns.
In intrusion detection, the FL-BERT model exhibits a high degree of precision and recall for Attack type 1 (Constant Attack) and Attack type 2 (Constant Offset Attack), with F1-scores of 0.95 and 1.00, respectively. This suggests that the model is highly effective at identifying and correctly classifying these types of attacks.
For Attack type 4 (Random Attack), the model shows a relatively lower precision of 0.65 but a high recall of 0.88, resulting in an F1-score of 0.75. This indicates that while the model correctly identifies a high proportion of this attack type (high recall), it also incorrectly classifies other types as this attack (lower precision).
The model’s performance on Attack type 8 (Random Offset Attack) and Attack type 16 (Eventual stop Attack) is mixed, with F1-scores of 0.55 and 0.86 respectively. The lower F1-score for Attack type 8 suggests that this type of attack is more challenging for the model to identify and classify correctly.
In conclusion, with its federated learning approach, the FL-BERT model demonstrates a promising capability in intrusion detection tasks, particularly in classifying different types of attacks. However, the varying performance across attack types suggests potential areas for further optimization and improvement.
FL-BERT’s advantages and disadvantages are similarly highlighted by evaluating using standard methods like Random Forest and SVM. Compared to FL-BERT, the much-reduced precision and do-not-forget values in these typical designs highlight how complicated the information is. Without the benefit of deep architectures, such models might not be able to grasp the subtle styles in VANET records, leading to more frequent misclassifications.
The imbalance of the records is a further important consideration. The training data structure is crucial in determining how a model makes predictions. The model’s knowledge may be skewed if certain assault types predominate in the dataset, making it either too aggressive or cautious. For instance, an excess of benign statistics may make the version prone to making warning errors, leading to the classification of capability risks as unimportant. In conclusion, although FL-BERT claims outstanding accuracy, understanding these nuances and continuously improving the version based only on these insights are crucial to minimizing misclassifications.
FL-BERT Privacy Concern
Implementing FL-BERT in a Software-defined Vehicular Ad-hoc Network (VANET) demonstrates the integration of advanced deep learning techniques with the intricate realm of vehicular communications. Vehicular Ad Hoc Networks (VANETs) are specifically engineered to introduce a novel era of intelligent transportation systems, wherein vehicles establish communication to ensure safety and optimize traffic flow. Nevertheless, the nature of statistics transmitted across these networks is sensitive. The continual sharing of information regarding every vehicle’s role, speed, route, and other factors raises privacy concerns, highlighting its importance.
Introducing FL-BERT, a customized iteration of the esteemed BERT model designed for federated learning. Federated Learning (FL) is a notable technology that enables models to acquire knowledge from several devices or servers without the requirement of centralizing data. The primary advantage of employing this technique lies in its ability to provide strict localization of raw recordings, which can be highly illuminating within the context of Vehicular Ad-Hoc Networks (VANETs). Instead of transmitting raw data, only updated model versions, such as gradients or weights, are returned to a central server. This implies that intricate details regarding specific vehicles, their routes, or communication patterns are never disclosed to any significant entity or possible individuals attempting to intercept such information.
Preserving anonymity is more appropriate when a central server collects updates from multiple clients to enhance the global model. The process of aggregating automatically provides a level of anonymization. By amalgamating updates from multiple sources, the task of reverse engineering or deducing specific details about an individual information source or user will become exceedingly challenging.
The periodic re-initialization of the version represents a widely implemented yet significant measure to enhance privacy. The gadget employs regularly resetting and disseminating the global version to all clients, preventing overfitting and excessive reliance on data from any specific user. The aforementioned behavior is a deterrent for potential attackers who may attempt to infer specific patterns from the actions of the version in question.
The significance of BERT, characterized by its transformer architecture, is equally crucial. Transformers exhibit a high level of proficiency in processing sequential data and have the capability to capture intricate patterns. In the context of Vehicular Ad-Hoc Networks (VANETs), this implies that the proposed model can effectively process vehicular data streams without the need for explicit feature engineering. This enhances the model’s accuracy and mitigates the potential risk of information leakage occasionally occurring with manually designed routines.
Nevertheless, it is important to note that no technology is completely impervious to vulnerabilities. Despite implementing federated studying, the persistence of vulnerabilities, like version inversion attacks, remains a concern. These attacks include malicious entities attempting to reconstruct raw information from version updates. To enhance the security measures of the gadget, incorporating additional techniques, such as differential privacy, would prove advantageous.
Using FL-BERT in VANETs is a significant step towards a future in which vehicular networks are more intelligent and prioritize each participant’s privacy. The results, which highlight the superior performance of FL-BERT compared to conventional models such as Random Forest and SVM, also underscore the potential of federated learning in these significant applications.
Conclusion
The need for robust security and privacy solutions is increasingly emphasized in the dynamic landscape of Software-defined Vehicular Ad Hoc Networks (VANETs). This study has made significant contributions in shedding light on a promising path forward, offering contemporary solutions that effectively negotiate the complex interplay between advanced machine learning and data privacy. The emergence of FL-BERT, a framework that combines Federated Learning with the BERT model, exemplifies the revolutionary potential of integrating decentralized learning with advanced text categorization algorithms. This integration not only effectively handles the urgent challenges of data privacy but also ensures that the effectiveness of risk detection is not affected.
The superiority of FL-BERT is further highlighted through our comparative evaluation with traditional machine learning models such as RF, SVM, LR, and KNN. FL-BERT consistently demonstrates impressive performance metrics, outperforming its more conventional counterparts, despite its intricate structure that prioritizes data privacy. The aforementioned findings demonstrate a significant shift in paradigms, implying that the future of secure VANETs may rely on using federated learning frameworks that exploit advanced device learning architectures.
Including the VeReMi dataset in our research has enhanced the severity of our empirical examination. The comprehensive scope of our study facilitated a thorough examination of our proposed methodologies, yielding valuable, enduring, illuminating insights. The dataset functioned as a testing ground, evaluating the effectiveness of our approaches against a wide range of security risks and, in doing so, identifying areas of strength and potential improvement.
Moreover, the observable scope facilitated by BERT’s contextual abilities reinforces the increasing agreement within the academic community about the revolutionary role of transformer-based models in threat detection. The adaptability and overall performance of the subject, particularly when compared to conventional models, provide a promising direction for future study efforts.
This study represents a significant advancement in pursuing secure and privacy-preserving Vehicular Ad Hoc Networks (VANETs). By advocating for implementing federated learning and leveraging advanced machine learning models, we present a comprehensive framework that enhances security and demonstrates a strong dedication to safeguarding data privacy. The insights and approaches presented in this study are expected to be crucial in shaping future advancements in VANETs. This study basically recommends implementing this solution to an SDN controller in VANET. [7]
Declarations
Competing interests
The authors declare that they have no competing interests.
Authors’ contributions
Work equally conducted by Shakil Ibne Ahsan and Co-authors.
Availability of data and materials
Not applicable.
Funding
Not applicable.
Acknowledgements
The genuine gratitude goes to the FET-Computer Science and Creative Technologies department, University of the West of England - UWE Bristol, for providing essential, necessary support to enrich this study. The authors would like to thank the reviewers for their valuable time.